### v1

# Pretraining Strategy for Neural Potentials

Abstract
We propose a mask pretraining method for Graph Neural Networks (GNNs) to improve their performance on fitting potential energy surfaces, particularly in water systems. GNNs are pretrained by recovering spatial information related to masked-out atoms from molecules, then transferred and finetuned on atomic forcefields. Through such pretraining, GNNs learn meaningful prior about structural and underlying physical information of molecule systems that are useful for downstream tasks. From comprehensive experiments and ablation studies, we show that the proposed method improves the accuracy and convergence speed compared to GNNs trained from scratch or using other pretraining techniques such as denoising. On the other hand, our pretraining method is suitable for both energy-centric and force-centric GNNs. This approach showcases its potential to enhance the performance and data efficiency of GNNs in fitting molecular force fields.
[…]
1 Introduction
Molecular dynamics plays a pivotal role in elucidating the dynamic behavior of molecules, providing essential insights into the temporal evolution of complex systems, such as understanding the conformational changes, interactions, and thermodynamic properties of molecules. It is thus important for fields ranging from drug discovery to materials science. The potential energy surface serves as the underlying landscape dictating the dynamics of a molecular system, and making an accurate representation of it is crucial for studying the dynamics of molecules through numerical simulations. Ab initio methods like Density Functional Theory (DFT) provide high accuracy by accounting for the electronic structure of atoms. Yet their high computational demands pose a significant challenge for efficient utilization, particularly in the context of large many-body systems. On the other hand, forces can be computed using empirical interatomic potentials tailored to specific environments, bypassing the electronic structures of the system. This approach significantly reduces the computational cost compared to ab initio methods. As there is a wide variety of interactions in molecular systems (bonded or non-bonded interactions), finding the appropriate functional forms can be challenging 1, 2. The difficulty in describing complex interatomic interaction presents a potential limitation for classical MD. Machine learning (ML) has emerged as a promising alternative for learning and fitting the potential energy surface with accuracy close to ab initio approaches 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14. The kernel-based regression model is one of the examples. Gradient-domain ML (GDML) employs a kernel-based method to fit the forcefield with the guarantee of energy conservation 15. Following GDML, Symmetric GDML (sGDML) extends GDML to incorporate space group symmetries and dynamic nonrigid symmetries 16. Gaussian approximation potentials (GAPs) is another notable ML method, which decomposes the energy as the sum of atomic-centered Gaussian basis function 17, 18 with hand-designed local environment descriptors 19, 20. The strong fitting capacity of the neural network has made it a suitable choice for fitting the potential energy surface. The initial efforts of building neural-network-based potential (neural potential) have leveraged tailored descriptors that capture the characteristics of local environment of atoms in the system 21, 7, 22, 23, 24, 25, where Behler-Parrinello Neural Networks (BPNN) 26, 27, 28 is one of the first examples. The recent advancements of deep neural networks, especially graph neural networks (GNNs), open up the possibility of learning the atomic representations directly from the raw atomic coordinates, which provides an alternative to the manually tailored atomic fingerprints 29, 30, 31, 32, 33, 34, 35. In GNNs, the interaction between atoms is modeled by message passing 36 between atoms within the local neighborhood. SchNet 33 introduces continuous convolution for aggregating messages across nodes (atoms) to produce a continuous energy surface. Various other works have developed different message-passing protocols based on physics and chemical insights 31, 32, 34, 35, 37. One notable variant is the class of equivariant GNNs. With tailored message-passing functions and update strategies, these neural networks are equivariant to group actions such as rotation and translation in 3D Euclidean groups. The exploitation of property of the symmetry has made equivariant GNNs not only more data-efficient and also more accurate 38, 39, 40, 41, 42, 43, 44.
While GNN-based neural potentials have provided a flexible and accurate end-to-end framework for learning and predicting the potential energy, their performance heavily relies on the amount of data. Collecting DFT-based data can be expensive or even infeasible for large and complex molecular systems. To improve the data efficiency of GNNs on molecular modeling tasks such as property prediction, various pretraining strategies have been investigated. One direction is pretraining GNNs to predict the masked attributes or context of the graph 45, 46, 47. Another direction is pretraining GNNs by contrasting the embeddings from the network given different views of the data 48, 49, 50, 51, 52, 53, 54, 47. Different views can be generated via randomly masking out some of the nodes and edges in the molecular graphs 48 or directly contrasting 3D representation with 2D representation of the molecular system 55, 56. It is worth noting that many of these methods have focused on improving the GNNs’ performance on 2D graphs, while for neural potential the input is usually 3D structures (coordinates of atoms) of molecular systems. Several recent works have proposed to pretrain the GNNs on 3D representation by denoising 57, 58, 59, 60. More specifically, the pretraining task is to predict the atom-wise Gaussian noise added to the atom coordinates under equilibrium state or non-equilibrium state. The denoising pretraining has been demonstrated to help improve models’ performance in predicting the properties of small molecules or single molecule systems. Previous works have viewed denoising as learning the pseudo-gradient (or force field) of the energy surface and an appropriately small perturbation (e.g. Gaussian noise with variance less than 0.1) can yield meaningful learning tasks for GNNs. However, identifying the appropriate noise scale can be difficult, such that it should not be trivially small nor excessively large. For the water system, the overall potential energy includes the contribution from intra-molecule interaction and inter-molecule interaction. This makes selecting an appropriate variance much trickier than single-molecule systems as small noise cannot yield significant enough perturbation to the intra-molecule part while large perturbation is too difficult for GNNs to recover the original geometric structures.
Figure 1: Framework of pretraining GNNs over masked water molecules. (a) Selecting water molecules with a certain ratio, spatial information of selected Hydrogen atoms (Blue) is masked to create a pretext task. (b) GNNs are pretrained to recover displacements between masked-out atoms and other atoms. Information on any atoms but masked-out ones is provided for GNNs to predict such displacements. (c) The pretrained GNNs are transferred and finetuned to predict the potential energy surface.
2 Methodology
2.1 Graph Neural Networks
Molecules are naturally suitable to be modeled as graphs, where atoms are depicted as nodes and their interactions as edges. Message Passing Neural Network (MPNN)36, has been a popular Graph Neural Network paradigm for its competitive performance on complexly structured graphs, which makes it apt for molecular property prediction. It often includes two phases during forward propagation: message passing and readout. We define the constructed graph $G$ that contains node features $x_{v}$ and edge features $e_{vw}$ encoded from atoms and bonds (interaction) between atoms. During the message passing stage, hidden states $h_{v}^{t}$ at each node are updated with an update function $f_{u}^{t}$ based on messages $m_{v}^{t+1}$ that are aggregated by a message function $f_{m}^{t}$.
$m_{v}^{(t+1)}=\sum_{j\in\mathcal{N}(i)}f_{m}^{(t)}(h_{v}^{(t)},h_{w}^{(t)},e_{vw})$ (1)
$h_{v}^{(t+1)}=f_{u}^{(t)}\left(h_{v}^{(t)},m_{v}^{(t+1)}\right)$ (2)
Then a readout function $R$ aggregates all node features to learn a graph-level representation.
$\hat{y}=R(\{h_{v}^{T}|v\in G\})$ (3)
Several research 61, 62, 45, focused on extending MPNNs to the field of molecular attribute predictions. Among them, a subset of equivariant and covariant MPNNs, EGNN63, GNS64, and ForceNet65 are used in our investigation as baseline models. EGNN extends equivariance to higher orders and shows competitive performance on molecular datasets such as QM9 66. When used to predict molecular dynamics, EGNN first predicts energy while maintaining rotational equivariant and energy conserving. The forces are then derived as the negative first derivatives of energy with respect to atomic positions. On the other hand, both GNS and ForceNet are based on encode-process-decode architecture. Generally, a GNN is used to encode and update node embeddings, which are then decoded and used to predict atomic forces. A Multilayer Perceptron (MLP) finally approximates the energy by aggregating decoded embeddings.
2.2 Pretraining by masking
Our proposed approach Hydrogen atom masking, similar to token masking in Masked Language Modeling 67, refers to deliberately masking one Hydrogen atom out of each selected water molecule with its spatial information removed from the input. The selection ratio is a hyperparameter that directly affects our framework’s performance. Analysis of masking ratios is further discussed in the supplementary information section 6.1.
Specifically, before masking, we calculate displacements $\mathbf{d}_{i}$ between atoms within the same molecule and append them to inputs along with indices of masked atoms, masked displacements and positions, atom types, and box sizes. Models will then predict the spatial information of the masked atoms and return their predicted displacements $\hat{\mathbf{d}}_{i}$ relative to other atoms in the same water molecules. The to-be minimized objective $\mathcal{L_{\text{masking}}}$ is the negative Cosine Similarity between the prediction $\hat{\mathbf{d}}_{i}$ and ground true label, $\mathbf{d}_{i}$:
$\hat{d_{i}}=\phi_{\theta}(V,\hat{X})$ (4)
$\mathcal{L_{\text{masking}}}=1-\frac{\sum_{i=1}^{n}\hat{d_{i}}\cdot d_{i}}{\sqrt{\sum_{i=1}^{n}\hat{d_{i}}^{2}}\cdot\sqrt{\sum_{i=1}^{n}d_{i}^{2}}},$ (5)
where $\phi_{\theta}$ denotes an GNN parametrized by $\theta$. $V$ and $X$ are the respective encoded atom information and Cartesian coordinates of input data samples.
The atom type we choose to mask out is based on the chemical properties of atoms in the water molecules. We empirically discovered that masking Oxygen atoms out causes instability during training and provides degraded results, which could partially be attributed to that an Oxygen atom weighs 18 times as a Hydrogen atom and exerts greater force on its neighbors. Consequently, we decided to mask out a hydrogen atom from selected water molecules and use it as the appropriate pretraining target.
On the other hand, the remaining Oxygen and Hydrogen atoms would have different impacts on the predictions of masked Hydrogen atoms’ coordinates due to their different atomic types. Thus, by reconstructing spatial information of missing atoms, GNNs are expected to learn from restoring the structural and physical information and the inherent structure of water systems. Moreover, the choice of using Cosine Similarity instead of $L1/L2$ distance to calculate the loss of displacement predictions is due to its scale invariance. It is more reasonable to have the GNNs learn the relative positioning of the atoms instead of the absolute values of the displacements, which may be of small variance under different MD simulation conditions such as changes in Temperature in the data generation process.
2.3 Pretraining by denoising
Denoising refers to predicting the artificial noise added to input atom coordinates in a self-supervised fashion. The noise $\xi\in\mathbb{R}^{N\times 3}$ is usually sampled from Gaussian $\mathcal{N}(0,\sigma I)$. Such noise is added to atom coordinates as perturbations. Taking noised atom coordinates, atom types, and box sizes as input, GNNs return the predicted noise $\hat{\mathbf{\xi}}$ added to the coordinates. The denoising objective is MSE Loss, aiming to capture and minimize the difference between predicted noise and true noise.
The denoising training objective is shown below.
$\hat{\xi_{i}}=\phi_{\theta}(V,\hat{X})$ (6)
$\mathcal{L_{\text{denoising}}}=\|\hat{\mathbf{\xi}}-\xi\|^{2}$ (7)
where $\phi_{\theta}$ denotes an GNN parametrized by $\theta$. $V$ and $X$ are the encoded atom information and noised Cartesian coordinates of input data samples.
The capacity to predict perturbation exerted over atom coordinates helps GNNs develop an understanding of reasonable positions and distances relative to each other. As analyzed in several prior works, denoising atomic coordinates is related to learning a pseudo-force field at the perturbed states 57, 68, 60, 69. For a given molecule with $N$ atoms and its corresponding coordinates $X$, the probability of this conformation ${X}$ is $p(X)\sim\exp{(-E(X))}$ following the Boltzmann distribution, with $E(X)$ being the potential energy at $X$. As shown in Vincent 70, the network $s_{\theta}$ the minimizes the following objective:
$\frac{1}{2}\mathbb{E}_{q_{\sigma}(\tilde{X}|X)p(X)}[||s_{\theta}(\tilde{X})-\nabla_{\tilde{X}}(\log q_{\sigma}(\tilde{X}|X)||_{2}^{2}]$ (8)
also satisfies $s_{\theta}(X)=\nabla_{X}\log q_{\sigma}(X)$, with $\tilde{X}=X+\xi$ and $\xi\sim\mathcal{N}(0,\sigma I)$. When $\sigma$ is small enough, $s_{\theta}$ serves as a good estimate of the score function of the true data distribution $\nabla_{X}\log p(X)$. Hence, the denoising pretraining using Gaussian noise with small variance is equivalent to estimating the score function or the negative gradient of the energy. However, this requires the variance to be small enough and we will show in the experiment that it is not a necessary condition to achieve good performance in practice.
Given denoising is a popular method to pretrain GNNs when approximating Molecular Dynamics, we also pretrained baseline models using denoising to offer a comparison, as indicated in section 3.4.
3 Experiment
3.1 Experiment Set up
The whole training pipeline is of two stages: pretraining and finetuning. In the pretraining stage, all models are firstly trained on a pretext task in a self-supervised way. Pretrained model weights are then transferred and finetuned with supervision during the finetuning stage. This transfer aims to leverage the knowledge learned about molecular systems during pretraining. Such knowledge embedded in the GNN parameters is expected to improve the performance of the GNN predicting force and energy labels with a strong prior. Specifically, we pretrain the baseline models with the aforementioned mask pretraining strategy in section 2.2. The overall architecture is illustrated in Figure1. This paradigm is model-agnostic and can easily be applied to different models.
The datasets used in this work are generated from classical molecular dynamics (MD) and density functional theory (DFT). The classical MD simulations are performed using OpenMM 71 with Tip3p forcefield 72. The systems considered are uniform single-atom or single-molecule systems with different simulation box sizes. Periodic Boundary Conditions (PBCs) are set in all directions, and a cut-off distance of 10 angstroms was used. NVT ensemble with a Nosé–Hoover chain thermostat 73, 74 is used for the simulation. We simulate each configuration of the molecular system to $50000$ steps with a time step size of $2.0$ femtoseconds and save the state of the system every 50 steps. Each configuration is generated by initializing particles in the system with random positions. A total of $10$ configurations are for generating the data. The DFT simulation data of water molecules are obtained from Cheng et al. 75, which uses the revised Perdew–Burke–Ernzerhof (RPBE) functional 76 with Van der Waals correction 77.
Two datasets, RPBE and Tip3p, are both split into a train-validation set and a test set with a 90% and 10% ratio, and the train-validation set is further split into a train set and a validation set following the same proportion. In our works, we reproduced three previously mentioned Graph Neural Networks, EGNN63, GNS64, and ForceNet65 as baseline models to evaluate our method’s performance.
During Pretraining, baseline models are trained on RPBE and Tip3p datasets for 100 and 25 epochs, respectively. The pretrained models are later finetuned over the same dataset. The finetuning training epochs are 300 epochs and 50 epochs respectively. The difference between pretraining and finetuning epochs mainly comes from the size of the dataset. RPBE has 7241 data samples. Meanwhile, in the Tip3p dataset, for a total of 10000 data points, each of them comes with 258 water molecules. Thus, GNNs should be trained for more epochs to make up for fewer data samples in the RPBE dataset. To better investigate the effectiveness of our masking pretraining strategy, we also trained GNNs from scratch without pretraining, and we kept the rest of the configurations the same, to compare with their pretrained counterparts. Training from scratch epochs are 300 epochs and 50 epochs respectively as a fair comparison.
We employ the AdamW optimizer 78 for both pretraining and finetuning. Learning rates during all stages of training are set to 1e-4 and eventually reduced to 1e-7, with a weight decay rate of 1e-4. To create an equal comparison, we apply the same batch sizes, 16 and 10, for RPBE and Tip3p datasets across training stages and models. Notably, we observed that EGNN is highly sensitive to larger batch sizes, as training would become fairly unstable. The cutoff radius for water systems is set to 3.4 angstroms.
3.2 Experiment Results
Datasets Models Pretrained Force RMSE Energy RMSE Force Improv Energy Improv Masking Ratio
RPBE EGNN 220.412 1597.309
EGNN 89.798 906.139 -58.87% -42.19% 0.25
GNS 214.970 1842.212
GNS 103.403 2217.729 -51.80% 20.31% 0.5
ForceNet - -
ForceNet 231.297 - - - 0.5
Tip3p EGNN 61.481 685.071
EGNN 39.291 241.853 -36.09% -64.70% 0.9
GNS 52.021 638.216
GNS 49.104 540.406 -5.61% -15.33% 0.5
ForceNet - -
ForceNet - - - - 0.5
Table 1: Quantitative force and energy performance analysis on RPBE and Tip3p datasets. Performance improvements are calculated between pretrained models and their trained-from-scratch counterparts. The units for force and energy RMSE are ($meV/$\mathrm{\SIUnitSymbolAngstrom}$$) and ($meV$).
Following the described training setup, we conducted experiments to investigate if masking pretraining would help the finetuning objective over energy and force prediction accuracy on molecule potential datasets. All the configurations are trained on four different seeds, and we report the averaged results. As shown in Table 1, we compared the RMSE of predictions of force and energy for our baseline models. The results show that our proposed masking pretraining strategy provides a significant improvement over both objectives across datasets. Most of the RMSE objectives are reduced by pretraining with our proposed masking pretext task. The average improvement of force and energy RMSE for EGNN and GNS is $38.09\%$ and $25.48\%$. Particularly, the pretrained EGNN model 63 gained more than $35\%$ performance improvement on all objectives. Masking pretraining also helps the GNS model 64 reduce $28.71\%$ force RMSE. This implies that our approach is model-agnostic (equivariant and covariant models) and can be generalized to both force-centric models and energy-centric models.
It is worth noting that the ForceNet model65 is hardly converging on both datasets without pretraining. Since ForceNet’s results are several orders of magnitude higher than other models’ results, we did not report out-of-scale results for clarity. However, we observe that finetuning ForceNet becomes more stable after pretraining in comparison to ForceNet trained from scratch. We also observe a similar trend that masked pretraining can effectively stabilize the training of other models. We argue that masked pretraining could facilitate more stable training and faster convergence. We elaborate on this point with training logs in the supplemental information section 6.2.
3.3 Comparison over models trained from scratch with additional epochs
Figure 2: Force and Energy performance comparison between models trained with regular and extended epochs: 300 epochs and 400 epochs on the RPBE dataset, and 50 epochs and 75 epochs on the Tip3p dataset. $\text{Improvement}=1-\frac{{RMSE}_{\text{(pretrained + regular epochs)}}}{{RMSE}_{\text{(train from scratch + regular/extended epochs)}}}$. A negative percentage indicates better performance after pretraining.
Extended experiments are also conducted to help demonstrate the efficacy and necessity of our two-stage training approach in comparison to simply extending the training duration without pretraining. In contrast to trained from scratch, pretrained models have yielded significant improvements on the RPBE dataset, as shown in figure 2. In the first row of figure 2, We compared the energy and force RMSE of our pretrained models to the models trained from scratch for either regular 300 or an extended 400 epochs. The results show that our training strategy of 100 epochs of pretraining followed by 300 epochs of finetuning still outperforms the elongated train-from-scratch models.
The performance gap in force RMSE between EGNN and GNS models trained on the RPBE dataset with and without pretraining sustains when compared to an extended training duration. It underscores the effectiveness of pretraining over merely increasing the training iterations. On Tip3p, the difference between the pretrained models and the model trained from scratch with 75 epochs is reduced on force RMSE. Energy-wise, a similar trend is observed. It is worth noting that the pretrained EGNN on the RPBE dataset achieves a higher energy RMSE improvement when compared with EGNN trained from scratch with 400 epochs instead of 300 epochs. We attribute it to the overfitting of the model to this dataset. This observation indicates that for models converging on smaller datasets given certain epochs, starting from a point where the model has learned meaningful molecule featurization (with pretraining), the network is less likely to converge to worse local minima during the final training phase.
Pretraining gives the models a better initialization and facilitates models to extract meaningful representation from the unlabelled data, resulting in more adept and competent models. Therefore, our findings suggests the integration of pretraining as an essential component in building machine learning potential training pipelines.
3.4 Comparison with Denoising
Figure 3: Force and Energy performance comparison between models pretrained with masking and denoising tasks
We provide the comparative performance metrics of our proposed masking and denoising methods, which is critical in evaluating the efficacy of computational models in predicting molecular properties. Figure 3 provided a direct comparison between these two methods across two datasets, for force and energy predictions. To create fair comparisons, we held masking pretrained models’ results from table 1 constant. We implemented denoising as described in Sec 2.3. The mean and standard deviation of the Gaussian Noise added are set to be 0 and 0.5 throughout the experiments.
We observe that in the RPBE dataset, the performance of denoising is similar and competitive to that of our method. We attribute this observation to the effectiveness of the denoising approach in single molecule and attribute label pairs datasets such as ISO17, and ANI-1x as demonstrated in recent research 57, 60. Containing on average 27 water molecules, data points in the RPBE dataset are comparatively smaller than those in the Tip3p dataset. Importantly, denoising has shown difficulties scaling to the Tip3p dataset. EGNN model under the masking condition has a energy RMSE of about 241.85 $meV$. The performance deteriorates significantly when denoising is applied, with poor convergence and RMSE rising to approximately 2442.75 $meV/$\mathrm{\SIUnitSymbolAngstrom}$$ by the end of the finetuning stage. Denoising also exhibits degraded performance on the Energy RMSE metric when pretraining using the GNS model on the Tip3p dataset. Both EGNN and GNS pretrained using masked pretraining have shown stable convergence during training without suffering from scalability and achieved superior performance.
4 Conclusion
We present a new method of pretraining for improving the performance of GNNs on forcefield learning. We demonstrate that our method could achieve better accuracy on both force and energy prediction. Moreover, it also facilitates more stable training and exhibits better convergence in comparison to models without pretraining or pretrained with other pretext tasks such as denoising. These results indicate the effectiveness of our method for further utilization of GNNs to facilitate the development of neural potentials. Scaling our method with larger and more advanced GNNs, training on larger-scale datasets and molecular systems, and extending our methods to auto-regressive models are promising directions for future work.
5 Data Availability Statements
The data and code that support the findings of this project can be found at: https://github.com/StevenZhang904/Masked_Pretrain.git.

### v2

@@
-Pretraining Strategy for Neural Potentials
+Masked Pretraining Strategy for Neural Potentials
 Abstract
-We propose a mask pretraining method for Graph Neural Networks (GNNs) to improve their performance on fitting potential energy surfaces, particularly in water systems.
-GNNs are pretrained by recovering spatial information related to masked-out atoms from molecules, then transferred and finetuned on atomic forcefields.
+We propose a masked pretraining method for Graph Neural Networks (GNNs) to improve their performance on fitting potential energy surfaces, particularly in water and small organic molecule systems.
+GNNs are pretrained by recovering spatial information of masked-out atoms from molecules selected with certain ratios, then transferred and finetuned on atomic forcefields.
 Through such pretraining, GNNs learn meaningful prior about structural and underlying physical information of molecule systems that are useful for downstream tasks.
-From comprehensive experiments and ablation studies, we show that the proposed method improves the accuracy and convergence speed compared to GNNs trained from scratch or using other pretraining techniques such as denoising.
-On the other hand, our pretraining method is suitable for both energy-centric and force-centric GNNs.
+With comprehensive experiments and ablation studies, we show that the proposed method improves both the accuracy and convergence speed of GNNs compared to their counterparts trained from scratch or with other pretraining techniques.
+In addition, our pretraining method is suitable for both energy-centric and force-centric GNNs.
 This approach showcases its potential to enhance the performance and data efficiency of GNNs in fitting molecular force fields.
@@
 Following GDML, Symmetric GDML (sGDML) extends GDML to incorporate space group symmetries and dynamic nonrigid symmetries 16.
-Gaussian approximation potentials (GAPs) is another notable ML method, which decomposes the energy as the sum of atomic-centered Gaussian basis function 17, 18 with hand-designed local environment descriptors 19, 20.
+Gaussian approximation potentials (GAPs) decompose the energy as the sum of atomic-centered Gaussian basis function 17, 18 with hand-designed local environment descriptors 19, 20.
 The strong fitting capacity of the neural network has made it a suitable choice for fitting the potential energy surface.
@@
 With tailored message-passing functions and update strategies, these neural networks are equivariant to group actions such as rotation and translation in 3D Euclidean groups.
-The exploitation of property of the symmetry has made equivariant GNNs not only more data-efficient and also more accurate 38, 39, 40, 41, 42, 43, 44.
+The exploitation of property of the symmetry has made equivariant GNNs not only more data-efficient and also more accurate 38, 39, 40, 41, 42, 43, 44, 45, 46.
 While GNN-based neural potentials have provided a flexible and accurate end-to-end framework for learning and predicting the potential energy, their performance heavily relies on the amount of data.
 Collecting DFT-based data can be expensive or even infeasible for large and complex molecular systems.
-To improve the data efficiency of GNNs on molecular modeling tasks such as property prediction, various pretraining strategies have been investigated.
-One direction is pretraining GNNs to predict the masked attributes or context of the graph 45, 46, 47.
-Another direction is pretraining GNNs by contrasting the embeddings from the network given different views of the data 48, 49, 50, 51, 52, 53, 54, 47.
-Different views can be generated via randomly masking out some of the nodes and edges in the molecular graphs 48 or directly contrasting 3D representation with 2D representation of the molecular system 55, 56.
+Pretraining and finetuning have been proven to be an effective popular paradigm that can facilitate better data efficiency 47, 48, 49.
+To specifically improve the data efficiency of GNNs on molecular modeling tasks such as property prediction, various pretraining strategies have been investigated.
+One direction is pretraining GNNs by predicting the masked attributes or context of graphs 50, 51, 52.
+Another direction is pretraining GNNs by contrasting the embeddings from the network given different views of the data 53, 54, 55, 56, 57, 58, 59, 52.
+Different views can be generated via randomly masking out some of the nodes and edges in the molecular graphs 53 or directly contrasting 3D representation with 2D representation of the molecular system 60, 61.
 It is worth noting that many of these methods have focused on improving the GNNs’ performance on 2D graphs, while for neural potential the input is usually 3D structures (coordinates of atoms) of molecular systems.
-Several recent works have proposed to pretrain the GNNs on 3D representation by denoising 57, 58, 59, 60.
-More specifically, the pretraining task is to predict the atom-wise Gaussian noise added to the atom coordinates under equilibrium state or non-equilibrium state.
+Several recent works have proposed to pretrain the GNNs on 3D representation by denoising 62, 63, 64, 65.
+More specifically, the pretraining task is to predict the atom-wise Gaussian noise added to the atom coordinates under the equilibrium state or non-equilibrium state.
 The denoising pretraining has been demonstrated to help improve models’ performance in predicting the properties of small molecules or single molecule systems.
-Previous works have viewed denoising as learning the pseudo-gradient (or force field) of the energy surface and an appropriately small perturbation (e.g.
+Previous works have viewed denoising as learning the pseudo-gradient (or force field) of the energy surface, as an appropriately small perturbation (e.g.
 Gaussian noise with variance less than 0.1) can yield meaningful learning tasks for GNNs.
@@
 For the water system, the overall potential energy includes the contribution from intra-molecule interaction and inter-molecule interaction.
-This makes selecting an appropriate variance much trickier than single-molecule systems as small noise cannot yield significant enough perturbation to the intra-molecule part while large perturbation is too difficult for GNNs to recover the original geometric structures.
-Figure 1: Framework of pretraining GNNs over masked water molecules.
-(a) Selecting water molecules with a certain ratio, spatial information of selected Hydrogen atoms (Blue) is masked to create a pretext task.
+This makes selecting an appropriate variance much trickier than single-molecule systems, given that small noise cannot yield significant enough perturbation to the intra-molecule part and large perturbation is too difficult for GNNs to recover the original geometric structures.
+In this work, we propose to pretrain GNNs with tasks of recovering the spatial information of selectively masked-out atoms, and we showcase that it’s robust and versatile for different molecular systems, from water systems to organic molecules.
+In addition, we observe that our proposed pretraining strategy yields consistent improvement on both non-equivariant and equivariant GNNs.
+Figure 1: Framework of masked pretraining GNNs on water molecule systems.
+(a) With water molecules selected on a certain ratio, spatial information of one Hydrogen atom (Blue) in each selected water molecule is masked to create a pretext task.
 (b) GNNs are pretrained to recover displacements between masked-out atoms and other atoms.
-Information on any atoms but masked-out ones is provided for GNNs to predict such displacements.
-(c) The pretrained GNNs are transferred and finetuned to predict the potential energy surface.
+Information of any atoms but the masked-out ones is provided for GNNs to predict such displacements.
+(c) The pretrained weights are transferred and finetuned to predict the potential energy surface.
 2 Methodology
@@
 During the message passing stage, hidden states $h_{v}^{t}$ at each node are updated with an update function $f_{u}^{t}$ based on messages $m_{v}^{t+1}$ that are aggregated by a message function $f_{m}^{t}$.
+Formally:
 $m_{v}^{(t+1)}=\sum_{j\in\mathcal{N}(i)}f_{m}^{(t)}(h_{v}^{(t)},h_{w}^{(t)},e_{vw})$ (1)
 $h_{v}^{(t+1)}=f_{u}^{(t)}\left(h_{v}^{(t)},m_{v}^{(t+1)}\right)$ (2)
-Then a readout function $R$ aggregates all node features to learn a graph-level representation.
+Then a readout function $R$ aggregates all node features to learn a graph-level representation:
 $\hat{y}=R(\{h_{v}^{T}|v\in G\})$ (3)
-Several research 61, 62, 45, focused on extending MPNNs to the field of molecular attribute predictions.
-Among them, a subset of equivariant and covariant MPNNs, EGNN63, GNS64, and ForceNet65 are used in our investigation as baseline models.
-EGNN extends equivariance to higher orders and shows competitive performance on molecular datasets such as QM9 66.
-When used to predict molecular dynamics, EGNN first predicts energy while maintaining rotational equivariant and energy conserving.
+Several research 66, 67, 50 focused on extending MPNNs to the field of molecular attribute predictions.
+Among them, a subset of equivariant and covariant MPNNs such as EGNN68, GNS69, and ForceNet70 are used in our investigation as baseline models.
+EGNN is an energy-centric model that extends equivariance to higher orders and shows competitive performance on molecular datasets such as QM9 71.
+When used to predict molecular dynamics, EGNN first predicts energy while maintaining rotational equivariant and energy conservation.
 The forces are then derived as the negative first derivatives of energy with respect to atomic positions.
-On the other hand, both GNS and ForceNet are based on encode-process-decode architecture.
-Generally, a GNN is used to encode and update node embeddings, which are then decoded and used to predict atomic forces.
-A Multilayer Perceptron (MLP) finally approximates the energy by aggregating decoded embeddings.
+On the other hand, both GNS and ForceNet are based on encode-process-decode architecture and are force-centric.
+Specifically, a GNN is used to encode and update node embeddings, which are then decoded and used to predict atomic forces; a Multilayer Perceptron (MLP) finally approximates the energy by aggregating decoded embeddings.
 2.2 Pretraining by masking
-Our proposed approach Hydrogen atom masking, similar to token masking in Masked Language Modeling 67, refers to deliberately masking one Hydrogen atom out of each selected water molecule with its spatial information removed from the input.
+Our proposed approach Hydrogen atom masking, similar to token masking in Masked Language Modeling 72, refers to deliberately masking one Hydrogen atom out of each selected water molecule with its spatial information removed from the input.
 The selection ratio is a hyperparameter that directly affects our framework’s performance.
-Analysis of masking ratios is further discussed in the supplementary information section 6.1.
+Analysis of the choice masking ratios is further discussed in the supplementary information Section 1.
 Specifically, before masking, we calculate displacements $\mathbf{d}_{i}$ between atoms within the same molecule and append them to inputs along with indices of masked atoms, masked displacements and positions, atom types, and box sizes.
-Models will then predict the spatial information of the masked atoms and return their predicted displacements $\hat{\mathbf{d}}_{i}$ relative to other atoms in the same water molecules.
+Models will then predict the spatial information of the masked atoms and return their predicted displacements $\hat{\mathbf{d}}_{i}$ relative to other atoms in the same molecule.
 The to-be minimized objective $\mathcal{L_{\text{masking}}}$ is the negative Cosine Similarity between the prediction $\hat{\mathbf{d}}_{i}$ and ground true label, $\mathbf{d}_{i}$:
@@
 where $\phi_{\theta}$ denotes an GNN parametrized by $\theta$. $V$ and $X$ are the respective encoded atom information and Cartesian coordinates of input data samples.
-The atom type we choose to mask out is based on the chemical properties of atoms in the water molecules.
-We empirically discovered that masking Oxygen atoms out causes instability during training and provides degraded results, which could partially be attributed to that an Oxygen atom weighs 18 times as a Hydrogen atom and exerts greater force on its neighbors.
-Consequently, we decided to mask out a hydrogen atom from selected water molecules and use it as the appropriate pretraining target.
-On the other hand, the remaining Oxygen and Hydrogen atoms would have different impacts on the predictions of masked Hydrogen atoms’ coordinates due to their different atomic types.
+The atom type we choose to mask out is based on the chemical properties of atoms in many molecules that are consist of hydrogens and other heavy atoms.
+On water system, we empirically discovered that masking Oxygen atoms out causes instability during training and yields degraded results, which could partially be attributed to that recovering Oxygen atom can be too challenging to provide meaningful training signal.
+Consequently, we decided to mask out a Hydrogen atom from selected molecules and use it as the appropriate pretraining target.
+On the other hand, the remaining atoms would have different impacts on the predictions of masked Hydrogen atoms’ coordinates due to their different atomic types.
 Thus, by reconstructing spatial information of missing atoms, GNNs are expected to learn from restoring the structural and physical information and the inherent structure of water systems.
@@
 2.3 Pretraining by Denoising
-Denoising refers to predicting the artificial noise added to input atom coordinates in a self-supervised fashion.
-The noise $\xi\in\mathbb{R}^{N\times 3}$ is usually sampled from Gaussian $\mathcal{N}(0,\sigma I)$.
-Such noise is added to atom coordinates as perturbations.
-Taking noised atom coordinates, atom types, and box sizes as input, GNNs return the predicted noise $\hat{\mathbf{\xi}}$ added to the coordinates.
-The denoising objective is MSE Loss, aiming to capture and minimize the difference between predicted noise and true noise.
+Here denoising refers to predicting the artificial noise added to the input atom coordinates in a self-supervised fashion.
+The noise $\xi\in\mathbb{R}^{N\times 3}$ is usually sampled from Gaussian $\mathcal{N}(0,\sigma I)$, and such noise is added to atom coordinates as perturbations.
+The capacity to predict perturbation exerted over atom coordinates helps GNNs understand the reasonable positioning of molecules and distances relative to each other.
+As analyzed in several prior works, denoising atomic coordinates is related to learning a pseudo-force field at the perturbed states 62, 73, 65, 74.
+For a given molecule with $N$ atoms and its corresponding coordinates $X$, the probability of this conformation ${X}$ is $p(X)\sim\exp{(-E(X))}$ following the Boltzmann distribution, with $E(X)$ being the potential energy at $X$.
+As shown in Vincent 75, the network $s_{\theta}$ the minimizes the following objective:
+$\frac{1}{2}\mathbb{E}_{q_{\sigma}(\tilde{X}|X)p(X)}[||s_{\theta}(\tilde{X})-\nabla_{\tilde{X}}(\log q_{\sigma}(\tilde{X}|X)||_{2}^{2}]$ (6)
+also satisfies $s_{\theta}(X)=\nabla_{X}\log q_{\sigma}(X)$, with $\tilde{X}=X+\xi$ and $\xi\sim\mathcal{N}(0,\sigma I)$.
+When $\sigma$ is small enough, $s_{\theta}$ serves as a good estimate of the score function of the true data distribution $\nabla_{X}\log p(X)$.
+Hence, the denoising pretraining using Gaussian noise with small variance is equivalent to estimating the score function or the negative gradient of the energy.
+Given denoising is a popular method to pretrain GNNs when approximating potential energy surface, we also pretrained baseline models using denoising to offer a comparison, as indicated in section 3.4.
+Specifically, taking noised atom coordinates, atom types, and box sizes as input, GNNs return the predicted noise $\hat{\mathbf{\xi}}$ added to the coordinates.
+We use the mean squared error (MSE) loss as the objective function of denoising pretraining to capture and minimize the difference between predicted and true noise.
 The denoising training objective is shown below.
@@
 where $\phi_{\theta}$ denotes an GNN parametrized by $\theta$. $V$ and $X$ are the encoded atom information and noised Cartesian coordinates of input data samples.
-The capacity to predict perturbation exerted over atom coordinates helps GNNs develop an understanding of reasonable positions and distances relative to each other.
-As analyzed in several prior works, denoising atomic coordinates is related to learning a pseudo-force field at the perturbed states 57, 68, 60, 69.
-For a given molecule with $N$ atoms and its corresponding coordinates $X$, the probability of this conformation ${X}$ is $p(X)\sim\exp{(-E(X))}$ following the Boltzmann distribution, with $E(X)$ being the potential energy at $X$.
-As shown in Vincent 70, the network $s_{\theta}$ the minimizes the following objective:
-$\frac{1}{2}\mathbb{E}_{q_{\sigma}(\tilde{X}|X)p(X)}[||s_{\theta}(\tilde{X})-\nabla_{\tilde{X}}(\log q_{\sigma}(\tilde{X}|X)||_{2}^{2}]$ (8)
-also satisfies $s_{\theta}(X)=\nabla_{X}\log q_{\sigma}(X)$, with $\tilde{X}=X+\xi$ and $\xi\sim\mathcal{N}(0,\sigma I)$.
-When $\sigma$ is small enough, $s_{\theta}$ serves as a good estimate of the score function of the true data distribution $\nabla_{X}\log p(X)$.
-Hence, the denoising pretraining using Gaussian noise with small variance is equivalent to estimating the score function or the negative gradient of the energy.
-However, this requires the variance to be small enough and we will show in the experiment that it is not a necessary condition to achieve good performance in practice.
-Given denoising is a popular method to pretrain GNNs when approximating Molecular Dynamics, we also pretrained baseline models using denoising to offer a comparison, as indicated in section 3.4.
 3 Experiment
-3.1 Experiment Set up
-The whole training pipeline is of two stages: pretraining and finetuning.
-In the pretraining stage, all models are firstly trained on a pretext task in a self-supervised way.
+3.1 Experiment Set up on Water Datasets
+The whole training pipeline is composed of two stages: pretraining and finetuning.
+In the pretraining stage, all baseline models are firstly trained on the masked pretraining strategy discussed in Section 2.2.
 Pretrained model weights are then transferred and finetuned with supervision during the finetuning stage.
@@
 Such knowledge embedded in the GNN parameters is expected to improve the performance of the GNN predicting force and energy labels with a strong prior.
-Specifically, we pretrain the baseline models with the aforementioned mask pretraining strategy in section 2.2.
 The overall architecture is illustrated in Figure1.
-This paradigm is model-agnostic and can easily be applied to different models.
-The datasets used in this work are generated from classical molecular dynamics (MD) and density functional theory (DFT).
-The classical MD simulations are performed using OpenMM 71 with Tip3p forcefield 72.
+This training framework is model-agnostic and can easily be applied to different models.
+The water datasets used in this work are generated from classical molecular dynamics (MD) and density functional theory (DFT).
+The classical MD simulations are performed using OpenMM 76 with Tip3p forcefield 77.
 The systems considered are uniform single-atom or single-molecule systems with different simulation box sizes.
 Periodic Boundary Conditions (PBCs) are set in all directions, and a cut-off distance of 10 angstroms was used.
-NVT ensemble with a Nosé–Hoover chain thermostat 73, 74 is used for the simulation.
+NVT ensemble with a Nosé–Hoover chain thermostat 78, 79 is used for the simulation.
 We simulate each configuration of the molecular system to $50000$ steps with a time step size of $2.0$ femtoseconds and save the state of the system every 50 steps.
@@
 A total of $10$ configurations are for generating the data.
-The DFT simulation data of water molecules are obtained from Cheng et al.
-75, which uses the revised Perdew–Burke–Ernzerhof (RPBE) functional 76 with Van der Waals correction 77.
-Two datasets, RPBE and Tip3p, are both split into a train-validation set and a test set with a 90% and 10% ratio, and the train-validation set is further split into a train set and a validation set following the same proportion.
-In our works, we reproduced three previously mentioned Graph Neural Networks, EGNN63, GNS64, and ForceNet65 as baseline models to evaluate our method’s performance.
-During Pretraining, baseline models are trained on RPBE and Tip3p datasets for 100 and 25 epochs, respectively.
+The DFT simulation data of water molecules are obtained from Morawietz et al.
+80, which uses the revised Perdew–Burke–Ernzerhof (RPBE) functional 81 with Van der Waals correction 82 and contains $7241$ structures in total.
+Two water datasets, RPBE and Tip3p, are each split into a train-validation set and a test set with a 90% and 10% ratio, and the train-validation set is further split into a train set and a validation set following the same proportion.
+In our works, we reproduced three previously mentioned Graph Neural Networks, EGNN68, GNS69, and ForceNet70 as baseline models to evaluate our method’s performance.
+GNNs are pretrained on RPBE and Tip3p datasets for 100 and 25 epochs, respectively.
 The pretrained models are later finetuned over the same dataset.
-The finetuning training epochs are 300 epochs and 50 epochs respectively.
-The difference between pretraining and finetuning epochs mainly comes from the size of the dataset.
+The finetuning epochs are 300 epochs and 50 epochs respectively.
+The difference between different sets of training epochs mainly comes from the size of the dataset.
 RPBE has 7241 data samples.
-Meanwhile, in the Tip3p dataset, for a total of 10000 data points, each of them comes with 258 water molecules.
-Thus, GNNs should be trained for more epochs to make up for fewer data samples in the RPBE dataset.
-To better investigate the effectiveness of our masking pretraining strategy, we also trained GNNs from scratch without pretraining, and we kept the rest of the configurations the same, to compare with their pretrained counterparts.
+Meanwhile, the Tip3p dataset has a total of 10000 data points, each of which comes with 258 water molecules.
+Thus, GNNs are trained for more epochs to make up for fewer data samples in the RPBE dataset.
+To better investigate the effectiveness of our masking pretraining strategy, we trained GNNs from scratch without pretraining as baseline models, and we kept the rest of the configurations the same to compare with their pretrained counterparts.
 Training from scratch epochs are 300 epochs and 50 epochs respectively as a fair comparison.
-We employ the AdamW optimizer 78 for both pretraining and finetuning.
+The specific computational cost of pretraining is discussed in Supplemental Information Section 3.
+We employ the AdamW optimizer 83 for both pretraining and finetuning.
 Learning rates during all stages of training are set to 1e-4 and eventually reduced to 1e-7, with a weight decay rate of 1e-4.
-To create an equal comparison, we apply the same batch sizes, 16 and 10, for RPBE and Tip3p datasets across training stages and models.
+We apply the same batch sizes, 16 and 10, for RPBE and Tip3p datasets across training stages and models.
 Notably, we observed that EGNN is highly sensitive to larger batch sizes, as training would become fairly unstable.
+Due to our limited computational resources and to create an equal comparison, all models’ hidden dimensions are set to 128.
 The cutoff radius for water systems is set to 3.4 Angstroms.
@@
 Datasets Models Pretrained Force RMSE Energy RMSE Force Improv Energy Improv Masking Ratio
-RPBE EGNN 220.412 1597.309
-EGNN 89.798 906.139 -58.87% -42.19% 0.25
-GNS 214.970 1842.212
-GNS 103.403 2217.729 -51.80% 20.31% 0.5
+RPBE EGNN 221.0 (81.3) 1570.1(660.7)
+EGNN 90.9(8.3) 907.6(96.1) -58.9% -42.2% 0.25
+GNS 214.2(74.8) 1842.2(616.4)
+GNS 103.2(1.3) 2216.4(76.9) -51.8% 20.3% 0.5
 ForceNet - -
-ForceNet 231.297 - - - 0.5
-Tip3p EGNN 61.481 685.071
-EGNN 39.291 241.853 -36.09% -64.70% 0.9
-GNS 52.021 638.216
-GNS 49.104 540.406 -5.61% -15.33% 0.5
+ForceNet 231.3(5.1) - - - 0.5
+Tip3p EGNN 61.5(4.0) 685.1(55.8)
+EGNN 39.3(2.0) 241.9(27.3) -36.1% -64.7% 0.9
+GNS 52.0(0.2) 638.2(11.6)
+GNS 49.1(0.3) 540.4(8.1) -5.6% -15.3% 0.5
 ForceNet - -
 ForceNet - - - - 0.5
-Table 1: Quantitative force and energy performance analysis on RPBE and Tip3p datasets.
-Performance improvements are calculated between pretrained models and their trained-from-scratch counterparts.
-The units for force and energy RMSE are ($meV/$\mathrm{\SIUnitSymbolAngstrom}$$) and ($meV$).
-Following the described training setup, we conducted experiments to investigate if masking pretraining would help the finetuning objective over energy and force prediction accuracy on molecule potential datasets.
-All the configurations are trained on four different seeds, and we report the averaged results.
-As shown in Table 1, we compared the RMSE of predictions of force and energy for our baseline models.
-The results show that our proposed masking pretraining strategy provides a significant improvement over both objectives across datasets.
+Table 1: Quantitative analysis of force and energy RMSE on RPBE and Tip3p datasets.
+Changes in RMSE are shown for pretrained models compared to their trained-from-scratch counterparts, with negative percentages indicating improvements.
+The units for force and energy RMSE are ($meV/$\mathrm{\SIUnitSymbolAngstrom}$$) and ($meV$), respectively.
+All results represent the average of four runs with different random seeds, with the corresponding standard errors reported in parentheses.
+Following the described training setup, we conducted experiments to investigate if masked pretraining would help the finetuning objective over energy and force prediction accuracy.
+All the configurations are trained four times with different random seeds, and we report the averaged results.
+As shown in Table 1, we compared the Root mean square error (RMSE) of predictions of force and energy.
+The results show that the masked pretraining strategy provides a significant improvement over both objectives across datasets.
 Most of the RMSE objectives are reduced by pretraining with our proposed masking pretext task.
 The average improvement of force and energy RMSE for EGNN and GNS is $38.09\%$ and $25.48\%$.
-Particularly, the pretrained EGNN model 63 gained more than $35\%$ performance improvement on all objectives.
-Masking pretraining also helps the GNS model 64 reduce $28.71\%$ force RMSE.
+Particularly, the pretrained EGNN model gained 47.48% and 53.45% of performance improvement on force and energy objectives respectively.
+Masking pretraining also helps the GNS model reduce $28.71\%$ of force RMSE.
 This implies that our approach is model-agnostic (equivariant and covariant models) and can be generalized to both force-centric models and energy-centric models.
-It is worth noting that the ForceNet model65 is hardly converging on both datasets without pretraining.
+It is worth noting that the ForceNet model is hardly converging on both datasets without pretraining.
 Since ForceNet’s results are several orders of magnitude higher than other models’ results, we did not report out-of-scale results for clarity.
 However, we observe that finetuning ForceNet becomes more stable after pretraining in comparison to ForceNet trained from scratch.
-We also observe a similar trend that masked pretraining can effectively stabilize the training of other models.
-We argue that masked pretraining could facilitate more stable training and faster convergence.
-We elaborate on this point with training logs in the supplemental information section 6.2.
+We also observe a similar trend that masked pretraining can effectively stabilize the training of other models, and we argue that masked pretraining could facilitate more stable training and faster convergence.
+We further demonstrate and elaborate on this point with training loss curves of the ForceNet model in the supplemental information Section 2.
 3.3 Comparison over models trained from scratch with additional epochs
@@
 In contrast to trained from scratch, pretrained models have yielded significant improvements on the RPBE dataset, as shown in figure 2.
-In the first row of figure 2, We compared the energy and force RMSE of our pretrained models to the models trained from scratch for either regular 300 or an extended 400 epochs.
+In the first row of figure 2, We compared the energy and force RMSE of our pretrained models to the models trained from scratch with either regular 300 or an extended 400 epochs.
 The results show that our training strategy of 100 epochs of pretraining followed by 300 epochs of finetuning still outperforms the elongated train-from-scratch models.
-The performance gap in force RMSE between EGNN and GNS models trained on the RPBE dataset with and without pretraining sustains when compared to an extended training duration.
+On the RPBE dataset, the performance gap in force RMSE between EGNN and GNS models trained with and without pretraining sustains, despite the additional finetuning time.
 It underscores the effectiveness of pretraining over merely increasing the training iterations.
-On Tip3p, the difference between the pretrained models and the model trained from scratch with 75 epochs is reduced on force RMSE.
-Energy-wise, a similar trend is observed.
+On Tip3p, the difference in force RMSE between the pretrained models and the model trained from scratch with extended 75 epochs is reduced, and a similar trend is observed energy-wise.
 It is worth noting that the pretrained EGNN on the RPBE dataset achieves a higher energy RMSE improvement when compared with EGNN trained from scratch with 400 epochs instead of 300 epochs.
@@
 Pretraining gives the models a better initialization and facilitates models to extract meaningful representation from the unlabelled data, resulting in more adept and competent models.
-Therefore, our findings suggests the integration of pretraining as an essential component in building machine learning potential training pipelines.
+Therefore, our findings suggest integrating masked pretraining into machine learning potential training pipelines is non-trivial and of great potential.
 3.4 Comparison with Denoising
-Figure 3: Force and Energy performance comparison between models pretrained with masking and denoising tasks
-We provide the comparative performance metrics of our proposed masking and denoising methods, which is critical in evaluating the efficacy of computational models in predicting molecular properties.
-Figure 3 provided a direct comparison between these two methods across two datasets, for force and energy predictions.
-To create fair comparisons, we held masking pretrained models’ results from table 1 constant.
+Figure 3: Force and Energy performance results from models pretrained with masking and denoising tasks on the RPBE and Tip3p datasets
+We provide the comparative performance metrics of our proposed masked and denoising methods, which is critical in evaluating the efficacy of different pretraining strategies in predicting molecular properties.
+Figure 3 provided a direct comparison between these two methods across two datasets for force and energy predictions.
+We held masking pretrained models’ results from table 1 constant.
 We implemented denoising as described in Sec 2.3.
@@
 We observe that in the RPBE dataset, the performance of denoising is similar and competitive to that of our method.
-We attribute this observation to the effectiveness of the denoising approach in single molecule and attribute label pairs datasets such as ISO17, and ANI-1x as demonstrated in recent research 57, 60.
-Containing on average 27 water molecules, data points in the RPBE dataset are comparatively smaller than those in the Tip3p dataset.
+We attribute this observation to the effectiveness of the denoising approach in single molecule and attribute label pairs datasets such as ISO17, and ANI-1x as demonstrated in recent research 62, 65.
+Containing on average 27 water molecules in each sample, the number of samples in the RPBE dataset is 77.17% less than that of the Tip3p dataset.
 Importantly, denoising has shown difficulties scaling to the Tip3p dataset.
-EGNN model under the masking condition has a energy RMSE of about 241.85 $meV$.
-The performance deteriorates significantly when denoising is applied, with poor convergence and RMSE rising to approximately 2442.75 $meV/$\mathrm{\SIUnitSymbolAngstrom}$$ by the end of the finetuning stage.
-Denoising also exhibits degraded performance on the Energy RMSE metric when pretraining using the GNS model on the Tip3p dataset.
+EGNN model with masked pretraining has an energy RMSE of about 241.85 $meV$.
+The performance deteriorates significantly when denoising is applied, with poor convergence and RMSE rising to 2442.75 $meV/$\mathrm{\SIUnitSymbolAngstrom}$$ by the end of the finetuning stage.
+Denoising also exhibits degraded performance on the Energy RMSE metric with the GNS model pretrained on the Tip3p dataset.
 Both EGNN and GNS pretrained using masked pretraining have shown stable convergence during training without suffering from scalability and achieved superior performance.
+3.5 Experiment on Revised MD17 Dataset
+We additionally evaluated our masked pretraining method with EGNN on the Revised MD17 dataset 84, which is a refined version of the MD17 dataset 85, 16.
+It contains molecular dynamics (MD) trajectories of 10 types of small organic molecules at DFT accuracy, each comes with 100,000 samples (except for Azobenzene which has 99,888 samples).
+Following the setup in Schütt et al.
+33, the training sample size is 50,000, with 1,000 samples for evaluation, and the rest for testing.
+For Pretraining, we combine all the training samples for each molecule as the pretraining dataset and further finetune the models on the individual molecules.
+Thus, the overall pretraining dataset is of size 500,000 for 10 molecules.
+After pretraining, the pretrained model checkpoint is subsequently finetuned on each molecule individually with the same data splits to ensure no data leakage.
+We adopt the number of pretraining epochs to 15 and finetuning epochs to 200 in addition to a batch size of 32.
+Following prior works, we use Adam Optimizer with a learning rate of 1e-4 and a minimum learning rate of 5e-5.
+Prior works 86, 87 regard force and energy weights as tunable hyperparameters during training.
+In this work, we set them to 40 and 15 respectively, with which we observed reasonable convergence of EGNN on the Revised MD17 dataset.
+We pretrained EGNN models using three masking ratios (0.25, 0.5, 0.75) and then fine-tuned them on the Uracil molecule as validation.
+Although all three models showed reasonable performance improvements, we transferred the model weights pretrained with a masking ratio of 0.5 and finetuned each molecule in the dataset on top of it for its best performance among all.
+Molecule Pretrain Force MAE Energy MAE Change (%)
+Force Energy
+Revised Uracile 22.0 11.8
+19.0 10.7 -13.6% -9.3%
+Revised Benzene 14.6 5.4
+7.7 6.2 -47.3% 14.8%
+Revised Naphthalene 14.6 30.3
+18.5 14.7 26.7% -51.5%
+Revised Aspirin 46.7 25.2
+29.5 16.0 -36.8% -36.5%
+Revised Malonaldehyde 32.9 14.4
+11.7 8.3 -64.4% -42.4%
+Revised Ethanol 16.8 8.8
+10.7 4.7 -36.3% -46.6%
+Revised Toluene 17.9 13.7
+12.1 10.2 -32.4% -25.5%
+Revised Salicylic Acid 27.5 25.9
+17.6 13.2 -36.0% -49.0%
+Table 2: Force MAE and Energy MAE performance change results for EGNN with or without masked pretraining on the Revised MD17 dataset.
+The units for force and energy MAE are ($meV/$\mathrm{\SIUnitSymbolAngstrom}$$) and ($meV$).
+We present the results for Force and Energy mean absolute error (MAE) in Table 2, showing an average reduction of 30.02% and 30.58%, respectively.
+These findings suggest that our method can effectively generalize to small organic molecules with diverse types of atoms that are heavier than the Hydrogen atom.
+A major advantage of our proposed strategy is that it can be trained in a diverse set of unlabelled configurations, after which the pretrained model can be transferred to different types of molecules.
+Additionally, initializing from the weights that have already learned general patterns from extensive data, the convergence of the pretrained model is also faster, allowing the model to quickly adapt to different molecular structures.
+Most importantly, this approach ensures consistent and reliable performance across different molecules, which is vital for robust and accurate potential energy approximation.
 4 Conclusion
-We present a new method of pretraining for improving the performance of GNNs on forcefield learning.
+We present a new method of masked pretraining for improving GNN’s performance on forcefield learning.
 We demonstrate that our method could achieve better accuracy on both force and energy prediction.
@@
 These results indicate the effectiveness of our method for further utilization of GNNs to facilitate the development of neural potentials.
-Scaling our method with larger and more advanced GNNs, training on larger-scale datasets and molecular systems, and extending our methods to auto-regressive models are promising directions for future work.
+Despite our method being currently limited to molecules that contain Hydrogen atoms, we believe that exploring masking other atoms or masking a combination of different types of atoms and scaling our method with larger and more advanced GNNs, are promising directions for future work.
+5 Supplementary Material
+Supplementary material includes the following sections:
+  • Section 1: Ablation study of the choice of masking ratios;
+  • Section 2: Further analysis of the ForceNet70 results mentioned in Sec 3.2;
+  • Section 3: Computational cost of masked pretraining;
 6 Data Availability Statements